OpenStudy (anonymous):

V=x(x-6)(x-8) Simplify the equation and write it in standard form.

OpenStudy (campbell_st):

well its V = x(x^2 - 14x + 48) or V = x^3 - 14x^2 + 48x not sure if thats what you need.

OpenStudy (anonymous):

can you explain what you did

OpenStudy (campbell_st):

I expanded the binomial product  this is multiplying the binomials to give the quadratic then its simply a case on multiplying every term by x